### v1

# Solution for Emotion Prediction Competition of Workshop on Emotionally and Culturally Intelligent AI

Abstract
This report provide a detailed description of the method that we explored and proposed in the WECIA Emotion Prediction Competition (EPC), which predicts a person’s emotion through an artistic work with a comment. The dataset of this competition is ArtELingo, designed to encourage work on diversity across languages and cultures. The dataset has two main challenges, namely modal imbalance problem and language-cultural differences problem. In order to address this issue, we propose a simple yet effective approach called single-multi modal with Emotion-Cultural specific prompt(ECSP), which focuses on using the single modal message to enhance the performance of multimodal models and a well-designed prompt to reduce cultural differences problem. To clarify, our approach contains two main blocks: (1)XLM-R[1] based unimodal model and X2-VLM[11] based multimodal model (2) Emotion-Cultural specific prompt. Our approach ranked first in the final test with a score of 0.627.
1 Introduction
Emotion prediction plays a critical role in text classification for tasks such as stance detection and lie detection[7]. As shown in Figure 1 , ArtElingo[5] dataset compares and contrasts annotations on WikiArt across language/culture. These differences are interesting and important. One might suggest using machine translation to translate English captions to many other languages, but this will miss much of the opportunity. So we should build human-compatible AI that is more aware of our emotional being is important for increasing the social acceptance of AI. In this competition, we used a language model and a vision-language pretrain model: XLM-R and X2-VLM as our baseline.
Consider Figure 1c, where an Arabic annotator assigned the image the label contentment, but the other two annotators used the label : sadness. Captions are useful for diving deeper into these differences. The sadness annotations mention death and disasters, in contrast with the contentment annotation that ends with : felling of satisfaction. Further, the Table 1 show the method provided by organizer. Obviously, using only one of these modes can’t achieve good results.The Vit can’t even surpass the accuracy of guessing, which is 1/9. The reason is already mentioned above. Therefore, we can find that language modal has more rich information and should be the main modal of training. The reason for the poor performance of BERT is also very obvious. BERT model didn’t consider multilingualism and cultural differences issues.
| Method | F1 | ACC |
| BERT | 0.56 | 0.68 |
| Vit | 0.01 | 0.00 |
To address these challenges, we propose single-multi modal with Emotion-Cultural specific prompt, as shown in Figure 2. Firstly, in order to improve the ability of mining information from language modal, we chose two different model. One from single modal and one from multimodal. By using the way of model ensemble, we hope different mining way will have a complementary effect. Secondly, we meticulously design a language prompt which based on the retrieval augmented method. Retrieval augmented method help us find the nearest sample in the training dataset. Thus, we can get a pseudo-label and construct our prompt.
We introduce single-multi modal with Emotion-Cultural specific prompt strategy, and its contributions can be summarized as follows:
We use the strategy of integrating single modal and multimodal models to enhance the mining of more informative language modalities.
2 Method
2.1 Base Model
In our work, we take XLM-R and X2-VLM as our base model.
2.1.1 XLM-R
XLM-R is a transformer-based [2] language model trained with the multilingual MLM objective on 100 languages, three languages in artelingo dataset included. In order to deal with multi-language issues, XLM-R proposed new methods for data processing and model optimization objectives. The former uses Sentence Piece with a unigram language model to build a shared sub-word vocabulary, and the latter introduces a supervised optimization objective of translation language modeling(TLM). In this competition, we directly added a linear layer to fine-tune the pre-trained model for the task of emotion prediction.
2.1.2 X2-VLM
Obviously, the task of competition is a multimodal emotion classification task. So we chose a vision-language pre-training model. X2-VLM, an all-in-one model with flexible modular architecture, in which further unify imgae-text pre-training and video-text pre-training in one model. So it’s able to learn unlimited visual concepts associated with diverse text descriptions. But the origin X2-VLM model can’t deal with multilingual issues. Therefore, we replaced the origin text encoder(BERT) with the XLM-R model mentioned before.
2.2 Emotion-Cultural specific prompt
In order to alleviate the modal imbalance[10] problem mentioned before and mine the potential information of text modality, we focus on designing a prompt [3]based on emotion and culture. We proposed the Emotion-Cultural specific prompt. It consists of two parts, namely simple prompt and prompt with retrieval augmentation[4].
2.2.1 Simple prompt
First, we build a simple prompt using the attributes in the dataset. Given a annotation, denoted as , it contains the following attributes: art_style, language, utterance, the first two of which provide some information about the image. Then we splice them as follows:
The art style of image is {art_style}. There is a comment from a {language} person. What emotions did he express? amusement, awe, contentment, excitement, anger, disgust, fear, sadness or something else,{utterance}.
Although, maybe the vision encoder and text encoder of X-VLM can learn the art style of image and the language information of text. But we found that adding this information can effectively improve the performance of the model. What’s more, adding the category information to the input text can help the model better understand the goals of the task.
2.2.2 Prompt with retrieval augmentation
To alleviate cultural differences issues, we use labels of samples who share the same language with the current sample to construct pseudo-labels and embed them into the simple prompt constructed before. The specific plan is as follows: Assume that the current sample is , and the embedding of its text and image are: , , which are obtained through text and image encoders (e.g. CLIP) respectively, and then splice them together to get [, ], and calculate the similarity with samples of the same language (cultural differences in the same language will be smaller), sort the similarities from large to small, retrieve the top-k samples(in this competition, we just chose the most similar sample which means the k equals one), and splice the labels of these samples as pseudo labels and embed them into prompt. The formulas to select pseudo label are as follows:
where we chose cosine similarity as the function, and is the hyperparameter, indicating the threshold of cosine similarity. The final prompt we constructed is as follows:
{Simple prompt}. {Input text}. The emotion this picture is most likely trying to express is {pseudo label}.
2.3 Test Time Augmentation
Considering that the effect of the image modality is not as good as the text modality, and to further improve the robustness of the multi-modal model X2-VLM, we use test-time augmentation enhancement technique in the inference stage which contains horizontal flipping, vertical flipping and random cropping of images, as shown in the figure 4. The following Experiments prove that it has a certain effect on the final results.
2.4 Discussion
Although prompt learning has a certain enhancement effect on multimodal classification, constructing a reasonable and robust prompt itself poses great challenges. Moreover, in this competition, the constructed prompt serves as generic input and cannot personalize the learning well. Addressing this issue remains an intriguing and open research challenge. Future work may build more efficient prompts to deal with the above problems.
3 Experiment
Dataset. Our method is fine-tuned on the pre-trained models of XLM-R and X2-VLM using official datasets. But We split the training and validation sets randomly instead of following the official way.
Metric. The evaluation metric for this task is the F1 Score, which is the harmonic mean of Precision and Recall.
Implementation Detail. The maximum number of text tokens used by the single-modal method is 90. The maximum number of text tokens for multi-modal model text is 100. The resolution of the image is 768×768. Our approach is founded on fine-tune principles and makes use of the pre-trained model made available on the official XLM-R 111 and X2-VLM website 222. Regarding hyperparameter settings, we configure the similarity threshold as
| Method | F1 | Acc |
| 0.56 | 0.68 | |
| 0.01 | 0.00 | |
| 0.613 | 0.725 | |
| +ecsp | 0.618 | 0.729 |
| X2-VLM | 0.619 | 0.730 |
| X2-VLM+ecsp | 0.622 | 0.735 |
| Model Ensemble | 0.627 | 0.741 |
Comparison Methods Result. Table 2 shows the F1 score performance, from which we can observe that: 1) Vit exhibits poor performance in the competition dataset, because there is a situation in the data set where a image corresponds to text with different emotions, as shown in figure 1. 2) X2-VLM outperforms baseline methods by achieving about 0.059 and 0.609 performance improvements respectively. This phenomenon exhibits that text modality plays a far more important role in emotion classification than image modality. Note that the single-modal model is the large version, but the multi-modal text encoder is the base version, but the effects of the two are very close, indicating that the image modality still has a positive effect on the results.
Ablation Study. To analyze the contribution of the prompt in our method, we conduct more ablation studies in this competition. The F1 score after adding different prompts is demonstrated in Table 3. From the table, we can clearly see that after adding prompt, the F1 value has a certain improvement, whether in single mode or multi-mode. At the same time, for the same model, after adding escp, the effect gain is greater than adding sp, which proves the effectiveness of the Emotion-Cultural specific prompt we designed. The results indicate that prompts contributes to improvement, and better prompts can lead to further enhancements.
| Method | F1 | Acc |
| 0.613 | 0.725 | |
| +sp | 0.614 | 0.728 |
| +pseudo-label | 0.612 | 0.724 |
| +ecsp | 0.618 | 0.729 |
| X2-VLM | 0.619 | 0.730 |
| X2-VLM+sp | 0.621 | 0.733 |
| X2-VLM+pseudo-label | 0.620 | 0.731 |
| X2-VLM+ecsp | 0.622 | 0.735 |
4 Conclusion
This report summarized our solution for the Emotion Recognition Challenge in the Workshop on Emotionally and Culturally Intelligent AI. Our approach was based on the Emotion-Cultural specific prompt and Test Time Augmentation, thereby effectively enhancing the overall performance of emotion recognition.